# A practical identifiability criterion leveraging weak-form parameter estimation

### v3

@@ Appendix A Test Function Selection
It is well known that the choice of test function has a significant impact on the performance of weak-form methods. There have been several efforts to wisely select test functions, including spectral matching [26] and quadrature error minimization [28]. And, more recently, a method has been proposed to construct test functions for accurate parameter estimation [55]. However, optimal selection of test functions remains an open problem. In this work, we choose the test function optimally, based on knowing the true parameter. This ensures that the focus of this work can remain on the impact of constructing weak-form Input/Output equations.
@@
Figure 9: The relative error in parameters $w$ when fitting model (9) to 400 observations using WENDy for polynomial test function choices of degree =[6,18]=[6,18]. The red star indicates the $12^{\text{th}}$ order polynomial used in this study.
Figure 10: The relative error in parameters $w$ when fitting model (9) to 400 observations using WENDy for polynomial test function radii of support $a=[0.45,0.65]$. The red star indicates the choice of $a=0.52$ used in this study.
@@
Figure 11: The relative error in parameters $\beta$ when fitting model (17) to 31 observations using WENDy for polynomial test function choices of degree =[10,26]=[10,26]. The red star indicates the $20^{\text{th}}$ order polynomial used in this study.
Figure 12: The relative error in parameters $\beta$ when fitting model (9) to 31 observations using WENDy for polynomial test function radii of support $a=[5,15]$. The red star indicates the choice of $a=11$ used in this study.
@@ Appendix C Structural Identifiability of Example Models
where $w_{1}=k_{21}V_{e}$, $w_{2}=k_{12}+k_{21}$, and $w_{3}=k_{12}+k_{21}+V_{e}$. In order to verify the structural identifiability from this input-output representation, we can consider Definition 2. However, for this to be a sufficient condition for the structural identifiability, the rank of the Wronskian of the differential monomials of the input-output representation must be checked [39]. Recently, [19] have developed a projection-based approach analogous to the differential elimination approach that takes into consideration conditions for the equivalence of multi and single experiment identifiability. We confirm both the form of equation (8) and that the model parameters are globally structurally identifiable using find_ioequations and assess_identifiability commands from the StructuralIdentifiability package in Julia [19]. While StructuralIdentifiability.jl also produces the above strong form input-output equation, we chose to use the Differential Algebra v4 package to generate equation (8) in order to perform the integration by parts procedure that is needed to generate the weak-form in this study.
While we cannot directly test the identifiability of the weak-form input-output equation (9), we can confirm the unique recoverablity of the parameters in the discretized WENDy equation $\mathbf{G}\mathbf{w}^{*}=\mathbf{b}$. In the case of no noise,
@@
Similarly, for the SIR example, we confirm both the form of equation (14) and that the model parameters are globally structurally identifiable using find_ioequations and assess_identifiability commands from the StructuralIdentifiability package in Julia [19].
@@ Appendix D Example 1 $(e,q)$-identifiability with lognormal noise
Figure 14: Using WENDy with equations (9) we find the SIR model to be generally practically identifiable for multiplicative noise between $0\%,20\%]$. (A.) The area in blue denotes where the model is (e,q)-identifiable, and the area in white denotes where the model is not (e,q)-identifiable. The (e,q)-identifiability was determined from 1,000 samples at each respective error level, e. The red star marks the $(10,20)$-identifiability criterion. (B.) The relative error determined from 1,000 samples at each respective error level, e. (C.) The proportion of estimated $95\%$ confidence intervals at $e=[0\%,20\%]\%$ that contain the true value of $w_{i}$.

### v4

@@ Appendix A Test Function Selection
It is well known that the choice of test function has a significant impact on the performance of weak-form methods. There have been several efforts to wisely select test functions, including spectral matching [MessengerBortz2021JComputPhys] and quadrature error minimization [BortzMessengerDukic2023BullMathBiol]. And, more recently, a method has been proposed to construct test functions for accurate parameter estimation [TranBortz2025arXiv250703206]. However, optimal selection of test functions remains an open problem. In this work, we choose the test function optimally, based on knowing the true parameter. This ensures that the focus of this work can remain on the impact of constructing weak-form Input/Output equations.
@@
Figure 9: The relative error in parameters $w$ when fitting model (9) to 400 observations using WENDy for polynomial test function choices of degree =[6,18]=[6,18]. The red star * indicates the $12^{\text{th}}$ order polynomial used in this study.
Figure 10: The relative error in parameters $w$ when fitting model (9) to 400 observations using WENDy for polynomial test function radii of support $a=[0.45,0.65]$. The red star * indicates the choice of $a=0.52$ used in this study.
@@
Figure 11: The relative error in parameters $\beta$ when fitting model (17) to 31 observations using WENDy for polynomial test function choices of degree =[10,26]=[10,26]. The red star * indicates the $20^{\text{th}}$ order polynomial used in this study.
Figure 12: The relative error in parameters $\beta$ when fitting model (9) to 31 observations using WENDy for polynomial test function radii of support $a=[5,15]$. The red star * indicates the choice of $a=11$ used in this study.
@@ Appendix C Structural Identifiability of Example Models
where $w_{1}=k_{21}V_{e}$, $w_{2}=k_{12}+k_{21}$, and $w_{3}=k_{12}+k_{21}+V_{e}$. In order to verify the structural identifiability from this input-output representation, we can consider Definition 2. However, for this to be a sufficient condition for the structural identifiability, the rank of the Wronskian of the differential monomials of the input-output representation must be checked [OvchinnikovPogudinThompson2023AAECC]. Recently, [DongGoodbrakeHarringtonEtAl2023SIAMJApplAlgebraGeometry] have developed a projection-based approach analogous to the differential elimination approach that takes into consideration conditions for the equivalence of multi and single experiment identifiability. We confirm both the form of equation (8) and that the model parameters are globally structurally identifiable using find_ioequations and assess_identifiability commands from the StructuralIdentifiability package in Julia [DongGoodbrakeHarringtonEtAl2023SIAMJApplAlgebraGeometry]. While StructuralIdentifiability.jl also produces the above strong form input-output equation, we chose to use the Differential Algebra v4 package to generate equation (8) in order to perform the integration by parts procedure that is needed to generate the weak-form in this study.
@@
Similarly, for the SIR example, we confirm both the form of equation (14) and that the model parameters are globally structurally identifiable using find_ioequations and assess_identifiability commands from the StructuralIdentifiability package in Julia [DongGoodbrakeHarringtonEtAl2023SIAMJApplAlgebraGeometry].
@@ Appendix D Example 1 $(e,q)$-identifiability with lognormal noise
Figure 14: Using WENDy with equations (9) we find the SIR model to be generally practically identifiable for multiplicative noise between $0\%,20\%]$. (A.) The area in blue denotes where the model is $(e,q)$-identifiable, and the area in white denotes where the model is not $(e,q)$-identifiable. The $(e,q)$-identifiability was determined from 1,000 simulations at each respective error level, e. The red star * marks the $(10,20)$-identifiability criterion. (B.) The relative error determined from 1,000 simulations at each respective error level, e. (C.) The proportion of estimated $95\%$ confidence intervals at $e=[0\%,20\%]\%$ that contain the true value of $w_{i}$.
@@
Appendix E Comparison to Profile Likelihood
We compare the $(e,q)$-identifiability for the blood-diffusion model given by equation (7), found using the weak-form equation error method, to the profile likelihood (PLA), found using an output error method, for two different choices of parameters $w$. As can be seen in Figure 15, we $(10,q)$-identifiability of the blood diffusion model parameters and profile likelihood identifiability with $10\%$ observational error are in agreement. Further, for observational error under $10\%$ we find agreement between the WENDy confidence intervals and the PLA confidence intervals, which are given in Table 2. In a different region of the parameter space, we find the blood-diffusion model is not $(e,q)$-identifiable for observational error greater $0.5\%$ and that this practical non-identifiability is driven by uncertainty in parameter $w_{3}$. We find agreement for these results using the profile likelihood method at a $3\%$ observational error level. It is important to note that for the figures and confidence intervals generated using the profile likelihood method, the true parameters were used as an initial guess for the first output error optimization step in order to avoid failure of convergence, which was a challenge for the output error method.
Figure 15: At $e=10\%$, the blood-diffusion model given by equation (7) with parameters $w=[30,6,12]$ is $(10,39)$-identifiable and practically identifiable by profile likelihood analysis when using 40 observations. (A.) Continuous dynamics of the $x_{1}$ compartment of model given by equation (7) with parameters $w=[30,6,12]$. (B.) The area in blue denotes where the model is $(e,q)$-identifiable, and the area in white denotes where the model is not $(e,q)$-identifiable. The $(e,q)$-identifiability was determined from 1,000 simulations at each respective error level, $e$. (C.) Profile likelihood (black line) for parameter $w_{1},$ $w_{2}$, $w_{3}$ where the red dashed line represents the $95\%$ confidence level threshold in output error objective function value.
$e$ Method $\text{CI}_{w_{1}}$ $\text{CI}_{w_{2}}$ $\text{CI}_{w_{3}}$
$1\%$ WENDy $[28.57,31.43]$ $[5.67,6.33]$ $[11.78,12.22]$
PLA $[26.63,34.96]$ $[5.09,7.34]$ $[11.07,13.31]$
$5\%$ WENDy $[26.57,33.43]$ $[5.21,6.79]$ $[11.40,12.59]$
PLA $[26.38,35.29]$ $[5.03,7.42]$ $[11.04,13.37]$
$10\%$ WENDy $[24.78,35.22]$ $[4.79,7.21]$ $[10.24,13.76]$
PLA $[26.24,35.58]$ $[4.99,7.53]$ $[11.01,13.44]$
Table 2: WENDy and Profile likelihood (PLA) $95\%$ confidence intervals for the blood-diffusion model given by equation (7) with true parameters $w=[30,6,12]$.
Figure 16: At $e=3\%$, the blood-diffusion model given by equation (7) with parameters $w=[0.1,2,2.1]$ is not $(e,q)$-identifiable and not practically identifiable by profile likelihood analysis when using 40 observations. (A.) Continuous dynamics of the $x_{1}$ compartment of the model given by equation (7) with parameters $w=[0.1,2,2.1]$. (B.) The area in blue denotes where the model is $(e,q)$-identifiable, and the area in white denotes where the model is not $(e,q)$-identifiable. The $(e,q)$-identifiability was determined from 1,000 simulations at each respective error level, $e$. (C.) Profile likelihood (black line) for parameter $w_{1},$ $w_{2}$, $w_{3}$ where the red dashed line represents the $95\%$ confidence level threshold in output error objective function value.
$e$ Method $\text{CI}_{w_{1}}$ $\text{CI}_{w_{2}}$ $\text{CI}_{w_{3}}$
$3\%$ WENDy $[0.044,0.19]$ $[0.57,2.79]$ [18.74,39.72][-18.74,39.72]
PLA [0.44,][-0.44,-] $[1.59,2.75]$ [,][-,-]
Table 3: Profile likelihood (PLA) $95\%$ confidence intervals for the blood-diffusion model given by equation (7) with true parameters $w=[0.1,2,2.1]$.